A square is transformed into a rectangle by increasing the length by 8 m and the width by 5 m, If the area of the resulting rectangle is 108 m^2, algebraically determine the length of each side of the original side.
Answer by mananth(16946)

A square is transformed into a rectangle by increasing the length by 8 m and the width by 5 m, If the area of the resulting rectangle is 108 m^2, algebraically determine the length of each side of the original side.

A square each side x
is transformed into a rectangle by increasing the length by 8 m = x+8
and the width by 5 m, =x+5
If the area of the resulting rectangle is 108 m^2,
(x+5)(x+8) = 108
x^2+13x+40=108
x^2+13x-68
x^2+17x-4x-68=0
x(x+17)-4(x+17)=0
(x+17)(x-4)=0
x=4 OR -17
ignore negative
So side of square = 4 m